Jack was 10 years older than Prisilla. together their ages totaled 166 years. what were their ages?
Answer by itscoolatul(2)

let
jack's age be x yrs
prisilla's age =x-10 yrs
given
x+x-10=166
2x-10=166
2x=176
x=88 yrs
jack's age =88yrs
prisilla's age =78 yrs